# Unleashing the Potential of Diffusion Models for Incomplete Data Imputation

### v1

@@ 5.2 Main Results
Next, we compare the results on the out-of-sample imputation tasks. Considering that some methods cannot be used for out-of-sample imputation, we reduce the number of compared methods. Additionally, for MOT [19], we employ the Round-Robin method proposed in its paper to adapt it for out-of-sample imputation tasks. Fig. 4 compares the MAEs and RMSEs in the out-of-sample imputation task. Comparing it with the results of in-sample imputation, we can easily observe that some methods exhibit significant performance differences between the two settings. For example, graph-based methods GRAPE and IGRM perform well in in-sample imputation, but their performance degrades significantly in out-of-sample imputation. IGRM even fails on all datasets in the out-of-sample imputation setting. In contrast, our DiffPuter demonstrates similarly excellent performance in both in-sample and out-of-sample imputation. This highlights DiffPuter’s superior performance and robust generalization capabilities.
Figure 4: MCAR: MAE and RMSE in out-of-sample imputation. RMSE on the News dataset are all missing since all imputation methods give rise to abnormal RMSE scores due to extreme values.
@@ 5.3 Ablation Studies
In this section, we conduct ablation studies to evaluate the efficacy of the individual designs of DiffPuter, e.g., iterative training, as well as the robustness of DiffPuter, e.g., performance under varying missing ratios.
Does iterative training improve Diffusion Imputation?
@@
In Fig. 6, we present the performance of DiffPuter’s imputation results from increasing EM iterations. Note that $k=0$ represents the imputation result of a randomly initialized denoising network, and $k=1$ represents the performance of a pure diffusion model without iterative refinement. It is clearly observed that a single diffusion imputation achieves only suboptimal performance, while DiffPuter’s performance steadily improves as the number of iterations increases. Additionally, we observe that DiffPuter does not require a large number of iterations to converge. In fact, $4$ to $5$ iterations are sufficient for DiffPuter to converge to a stable and satisfying state.
Figure 5: Performance at different iterations.
Figure 6: Impacts of missing ratios.
Performance under varying missing ratios.
Finally, we investigate if the performance of DiffPuter will be highly impacted by the missing ratios compared with other imputation methods. In Fig. 6, we compare DiffPuter’s MAE score on the Beijing dataset’s continuous features with other SOTA imputation methods under missing ratio $r=0.3,0.5,$ and $0.7$. We find that the most competitive baseline method, Remasker, performs well under low missing ratios. However, when the missing ratio becomes large, its performance drops significantly, being surpassed by all other methods. Our DiffPuter exhibits the most robust performance across increasing missing ratios. Even with a missing ratio of $0.7$, it maintains very good imputation performance.
@@ 6 Conclusions
Appendix A Proofs
A.1 Proof for Theorem 1
@@
$\begin{split}p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})&=\int p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}_{t},\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs}){\rm d}\mathbf{x}_{t}\\ &={\mathbb{E}}_{p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})}p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}_{t},\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})\\ &\approx{\mathbb{E}}_{p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})}p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}_{t})\\ &\approx p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\tilde{\mathbf{x}}_{t}),\\ \end{split}$ (11)
where $\tilde{\mathbf{x}}_{t}$ is a random sample from $p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}^{\rm obs}_{0})$. The first $\approx$’ holds because when $\Delta t\rightarrow 0$, $\mathbf{x}^{\rm mis}_{t-\Delta t}$ is almost predictable via $\mathbf{x}_{t}$ without $\mathbf{x}_{0}^{\rm obs}$. The second $\approx$’ is due to Monte Carlo estimation. Therefore, when $\tilde{\mathbf{x}}_{t}\sim p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$, $p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$ is approximately tractable via $p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\tilde{\mathbf{x}}_{t})$.
@@
Appendix B Further explanation of the Diffusion Model
The diffusion model we adopt in Section 4.1 is actually a simplified version of the Variance-Exploding SDE proposed in  [27].
Note that  [27] has provided a unified formulation via the Stochastic Differential Equation (SDE) and defines the forward process of Diffusion as
@@ Appendix B Further explanation of the Diffusion Model
Variance Exploding (VE) implements the perturbation kernel Eq. 15 by setting $s(t)=1$, indicating that the noise is directly added to the data rather than weighted mixing. Therefore, The noise variance (the noise level) is totally decided by $\sigma(t)$. The diffusion model used in DiffPuter belongs to VE-SDE, but we use linear noise level (i.e., $\sigma(t)=t$) rather than $\sigma(t)=\sqrt{t}$ in the vanilla VE-SDE [27]. When $s(t)=1$, the perturbation kernels become:
@@
Appendix C Experimental Details
C.1 Configurations.
@@ Appendix C Experimental Details
  • Software: CUDA 12.2, Python 3.9.16, PyTorch [22] 1.12.1
C.2 Datasets
@@ C.3 Missing mechanisms
The code for generating masks according to the three missing mechanisms is also provided in the supplementary.
C.4 Implementations and hyperparameters
We use a fixed set of hyperparameters, which will save significant efforts in hyperparameter-tuning when applying DiffPuter to more datasets. For the diffusion model, we set the maximum time $T=80$, the noise level $\sigma(t)=t$, which is linear to $t$. The score/denoising neural network $\bm{\epsilon}(\mathbf{x}_{t},t)$ is implemented as a $5$-layer MLP with hidden dimension $1024$. $t$ is transformed to sinusoidal timestep embeddings and then added to $\mathbf{x}_{t}$, which is subsequently passed to the denoising function. When using the learned diffusion model for imputation, we set the number of discrete steps $M=50$ and the number of sampling times per data sample $N=10$. DiffPuter is implemented with Pytorch, and optimized using Adam [11] optimizer with a learning rate of $1\times 10^{-4}$.
@@ C.4 Implementations and hyperparameters
We use the same architecture of denoising neural network as in two recent tabular diffusion models for tabular data synthesis [13, 32]. The denoising MLP takes the current time step $t$ and the feature vector $\mathbf{x}_{t}\in{\mathbb{R}}^{1\times d}$ as input. First, $\mathbf{x}_{t}$ is fed into a linear projection layer that converts the vector dimension to be $d_{\rm hidden}=1024$:
@@
Then, following the practice in TabDDPM [13], the sinusoidal timestep embeddings $\mathbf{t}_{\rm emb}\in{\mathbb{R}}^{1\times d_{\rm hidden}}$ is added to $\mathbf{h}_{0}$ to obtain the input vector $\mathbf{h}_{\rm hidden}$:

### v2

@@ 5.2 Main Results (Mask Completely at Random)
Next, we compare the performance in the out-of-sample imputation tasks. Noting that some imputation methods are specifically designed for in-sample imputation and cannot be applied to the out-of-sample setting, the number of baselines in this task is significantly reduced. Table 6 in Appendix E.1 compares the MAEs and RMSEs in the OOS imputation task. Comparing it with the results of in-sample imputation, we can easily observe that some methods exhibit significant performance differences between the two settings. For example, graph-based methods GRAPE and IGRM perform well in in-sample imputation, but their performance degrades significantly in out-of-sample imputation. IGRM even fails on all datasets in the out-of-sample imputation setting. In contrast, our DiffPuter demonstrates excellent performance in both in-sample and out-of-sample imputation. This highlights DiffPuter’s superior performance and robust generalization capabilities.
Figure 3: Impacts of the number of EM iterations. DiffPuter’s performance steadily improves as the number of EM interactions increases.
Figure 4: Impacts of the number of sampled imputations per iteration. A very small $N$ leads to poor performance and large variance.
@@ 5.3 Ablation Studies
Impacts of the number of EM iterations. In Fig. 4, we present the performance of DiffPuter’s imputation results from increasing EM iterations. Note that $k=0$ represents the imputation result of a randomly initialized denoising network, and $k=1$ represents the performance of a pure diffusion model without iterative refinement. It is clearly observed that a single diffusion imputation achieves only suboptimal performance, while DiffPuter’s performance steadily improves as the number of iterations increases. Additionally, we observe that DiffPuter does not require a large number of iterations to converge. In fact, $4$ to $5$ iterations are sufficient for DiffPuter to converge to a stable and satisfying state.
Impacts of the number of sampled imputations per iteration. To obtain the expected value of missing entries, we have to sample a number of examples from the conditional distribution (via the reverse diffusion process). Therefore, the number $N$ should have a huge impact on DiffPuter’s performance. In Figure 4, we study the impacts of the number of samples on the imputation results. As expected, a very small $N$ leads to poor performance and large variance, while increasing the sampling number consistently improves the performance and reduces the variance. An optimal and stable performance can be achieved at $N\geq 10$. Since the sampling time is linear w.r.t. to the sample number, we use $N=10$ by default.
Figure 5: Impacts of sampling steps on the training time and imputation performance (MAE). Reducing diffusion sampling steps greatly reduces the training time at the cost of a slight performance drop.
Figure 6: Impacts of extremely large missing ratios. The performance of DiffPuter is upper-bounded by the initialized missing values via mean imputation
Impacts of the number of sampling steps in diffusion. Another hyperparameter impacting the performance and training speed is the number of sampling steps $M$ in the reverse process of diffusion. By default, DiffPuter set $M=50$, and we study the impact of reducing $M$ in  Figure 6. As demonstrated, reducing the diffusion sampling steps can greatly reduce the training time at the cost of only a slight performance drop. More detailed, reducing the sampling step from 50 to 20 reduces about $25\%$ of training time cost, with only $3\%$ performance drop. Therefore, when time costs are high, and precision requirements are not stringent, the training speed can be further increased by reducing the sampling steps.
Impacts of large missing ratios. In  Figure 6, we study the performance of DiffPuter with missing ratio from $30\%$ to $99\%$ (observed ratio from $70\%$ to $1\%$). Consistent with the intuition, when the observed ratio decreases (i.e., missing ratio increases), there is a significant increase in the imputation MAE score, indicating that the imputation performance becomes increasingly worse. However, we can observe a clear performance lower bound when the missing ratio approaches $99\%$, which is close to the performance of directly imputing using the mean of observed values. This is because, in DiffPuter’s first E-step, each column’s missing values are always initialized as the mean of the observed ones. Therefore, DiffPuter can still give a reasonable imputation result even if the missing ratio is extremely large.
Table 2: Comparison of the real training time on an Nvidia RTX 4090 GPU. DiffPuter has a similar training cost to SOTA methods, yet brings from $8\%$ to $25\%$ performance improvement.
Datasets MOT TDM GRAPE IGRM HyperImpute Remasker DiffPuter
Time: California $446.47s$ $216.91s$ $635.7s$ $1267.5s$ $1277.3s$ $1320.1s$ $1927.2s$
Time: Adult $396.68s$ $514.56s$ $2347.1s$ $3865.1s$ $1806.9s$ $1902.4s$ $2142.9s$
DiffPuter’s perf. improve. $21.47\%$ $21.37\%$ $25.94\%$ $20.97\%$ $8.44\%$ $10.65\%$
Comparison of Training Time. Intuitively, the training of DiffPuter would be very time-consuming, as it requires iteratively alternating between the training and sampling processes of the diffusion model. However, in practice, DiffPuter’s training efficiency is significantly improved through the following design: 1) light-weighted MLP-architectured denoising function; 2) reduced number of sampling steps ($M=50\ll 1000$ in traditional diffusion methods). To demonstrate this, we compare the training time of DiffPuter with other SOTA imputation methods. As demonstrated in Table 2, DiffPuter has similar training cost (similar scale) to these representative SOTA methods, yet brings from 8% to 25% performance improvement, which is acceptable.
Table 3: Effects of combining EM with other Deep Generative Models.
Methods Adult Default Shoppers News
MIWAE $0.5763$ $0.5194$ $0.5047$ $0.6349$
HIWAE $0.6155$ $0.3989$ $0.4707$ $0.5032$
VAEM $0.5568$ $0.4292$ $0.4626$ $0.5204$
HH-VAEM $0.5673$ $0.4589$
EM + MIWAE $0.5661$ $0.4993$ $0.4554$ $0.4171$
EM + HIWAE $0.5974$ $0.4314$ $0.4961$ $0.5222$
EM + VAEM $0.5492$ $0.4121$ $0.4362$ $0.5045$
EM + HH-VAEM $0.5402$ $0.4262$
DiffPuter $\bf{{\color[rgb]{0.9,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.2,0.2}0.3425}}$ $\bf{{\color[rgb]{0.9,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.2,0.2}0.2661}}$ $\bf{{\color[rgb]{0.9,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.2,0.2}0.3485}}$ $\bf{{\color[rgb]{0.9,0.2,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.9,0.2,0.2}0.2855}}$
Incorporating EM with other DGMs. Finally, we try to combine EM algorithms with other deep generative models that are easy to perform conditional sampling, including MIWAE (Mattei & Frellsen, 2019), HIWAE (Nazabal et al., 2020a), VAEM (Ma et al., 2020) and HH-VAEM (Peis et al., 2022a). Based on Table 3, our experimental results demonstrate that combining EM with other Deep Generative Models leads to performance improvements. However, it’s noteworthy that despite these improvements, our proposed DiffPuter still outperforms all combinations, achieving superior results, where lower scores indicate better performance. This demonstrates the effectiveness and robustness of our approach compared to existing methods, even when they are enhanced through combination with EM.
@@ 6 Conclusions
Acknowledgments
This work is supported in part by NSF under grants III-2106758, and POSE-2346158.
Appendix A Definition of Symbols
In Table 4, we explain the definition of every symbol used in this paper.
Table 4: Explanations of the symbols used in this paper.
Symbol Explanation
$d$ the number of columns of tabular data
$\mathbf{x}\in{\mathbb{R}}^{d}$ random variable, representing a row of the tabular data
$p(\mathbf{x})=p_{\rm data}(\mathbf{x})$ the distribution of the complete data
$\mathbf{m}\in\{0,1\}^{d}$ binary mask vector
$\mathbf{x}^{\rm obs}$ the observed part of $\mathbf{x}$, indicated by $m_{k}=0$
$\mathbf{x}^{\rm mis}$ the missing part of $\mathbf{x}$, indicated by $m_{k}=1$
$t$ the timestep of diffusion process
$\sigma(t)=t$ the noise level at timestep $t$
$T$ the maximum timestep
$\varepsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ standard Gaussian noise
$\mathbf{x}_{0}=\mathbf{x}$ unforwarded $\mathbf{x}$ (at timestep 00)
$\mathbf{x}^{\rm obs}_{0}=\mathbf{x}^{\rm obs}$ observed part of $\mathbf{x}$
$\mathbf{x}^{\rm mis}_{0}=\mathbf{x}^{\rm mis}$ missing part of $\mathbf{x}$
$\mathbf{x}_{t}=\mathbf{x}_{0}+\sigma(t)\varepsilon$ forwarded $\mathbf{x}$ at timestep $t$
$\mathbf{x}^{\rm obs}_{t}=\mathbf{x}^{\rm obs}_{0}+\sigma(t)\varepsilon$ observed part of $\mathbf{x}_{t}$
$\mathbf{x}^{\rm mis}_{t}=\mathbf{x}^{\rm mis}_{0}+\sigma(t)\varepsilon$ missing part of $\mathbf{x}_{t}$
$\mathbf{\theta}$ diffusion model’s parameters
$p_{\mathbf{\theta}}(\mathbf{x})$ density of data induced by the diffusion model
$\tilde{\mathbf{x}}_{T}\sim\pi(\mathbf{x})={\mathcal{N}}(\mathbf{0},\sigma^{2}(T)\mathbf{I})$ sampled random noise for the reverse process
$\tilde{\mathbf{x}}_{T}^{\rm obs}$ observed part of $\tilde{\mathbf{x}}_{T}$
$\tilde{\mathbf{x}}_{T}^{\rm mis}$ missing part of $\tilde{\mathbf{x}}_{T}$
$\tilde{\mathbf{x}}_{t}$ sampled data at timestep $t$ in the reverse process
$\tilde{\mathbf{x}}_{t}^{\rm obs}$ observed part of $\tilde{\mathbf{x}}_{t}$
$\tilde{\mathbf{x}}_{t}^{\rm mis}$ missing part of $\tilde{\mathbf{x}}_{t}$
$\hat{\mathbf{x}}_{0}^{\rm obs}$ The exact ground-truth value(s) of the observed part
Appendix B Proofs
B.1 Proof for Theorem 1
@@
$\begin{split}p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})&=\int p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}_{t},\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs}){\rm d}\mathbf{x}_{t}\\ &={\mathbb{E}}_{p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})}p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}_{t},\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})\\ &\approx{\mathbb{E}}_{p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})}p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}_{t})\\ &=p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\tilde{\mathbf{x}}_{t}),\\ \end{split}$ (11)
where $\tilde{\mathbf{x}}_{t}$ is a random sample from $p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}^{\rm obs}_{0})$. The first $\approx$’ holds because when $\Delta t\rightarrow 0$, $\mathbf{x}^{\rm mis}_{t-\Delta t}$ is almost predictable via $\mathbf{x}_{t}$ without $\mathbf{x}_{0}^{\rm obs}$. Therefore, when $\tilde{\mathbf{x}}_{t}\sim p_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$, $p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\mathbf{x}^{\rm obs}=\hat{\mathbf{x}}_{0}^{\rm obs})$ is approximately tractable via $p_{\bm{\theta}}(\mathbf{x}^{\rm mis}_{t-\Delta t}|\tilde{\mathbf{x}}_{t})$.
@@
Appendix C Further explanation of the Diffusion Model
The diffusion model we adopt in Section 4.1 is actually a simplified version of the Variance-Exploding SDE proposed in  (Song et al., 2021b).
Note that  (Song et al., 2021b) has provided a unified formulation via the Stochastic Differential Equation (SDE) and defines the forward process of Diffusion as
@@ Appendix C Further explanation of the Diffusion Model
Variance Exploding (VE) implements the perturbation kernel Eq. 15 by setting $s(t)=1$, indicating that the noise is directly added to the data rather than weighted mixing. Therefore, The noise variance (the noise level) is totally decided by $\sigma(t)$. The diffusion model used in DiffPuter belongs to VE-SDE, but we use linear noise level (i.e., $\sigma(t)=t$) rather than $\sigma(t)=\sqrt{t}$ in the vanilla VE-SDE (Song et al., 2021b). When $s(t)=1$, the perturbation kernels become:
@@
Appendix D Experimental Details
D.1 Configurations.
@@ Appendix D Experimental Details
  • Software: CUDA 12.2, Python 3.9.16, PyTorch (Paszke et al., 2019) 1.12.1
D.2 Datasets
We use ten real-world datasets of varying scales, and all of them are available at Kaggle [^2] https://www.kaggle.com/ or the UCI Machine Learning repository [^3] https://archive.ics.uci.edu/. We consider five datasets of only continuous features: California [^4] https://www.kaggle.com/datasets/camnugent/california-housing-prices, Letter [^5] https://archive.ics.uci.edu/dataset/59/letter+recognition, Gestur [^6] https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation, Magic [^7] https://archive.ics.uci.edu/dataset/159/magic+gamma+telescope, and Bean [^8] https://archive.ics.uci.edu/dataset/602/dry+bean+dataset, and five datasets of both continuous and discrete features: Adult [^9] https://archive.ics.uci.edu/dataset/2/adult, Default [^10] https://archive.ics.uci.edu/dataset/350/default+of+credit+card+clients, Shoppers [^11] https://archive.ics.uci.edu/dataset/468/online+shoppers+purchasing+intention+dataset, and News [^12] https://archive.ics.uci.edu/dataset/332/online+news+popularity. The statistics of these datasets are presented in Table 5.
@@ D.3 Missing mechanisms
We follow the methods proposed in Zhao et al. (2023) to implement MAR and MNAR.
  • "For MAR, we first sample a subset of features (columns in $X$) that will not contain missing values, and then we use a logistic model with these non-missing columns as input to determine the missing values of the remaining columns, and we employ line search of the bias term to get the desired proportion of missing values."
  • For MNAR, we use the first approach proposed in [4], "Using a logistic model with the input masked by MCAR". Specifically, similar to the MAR setting, we first divide the data columns into two groups, with one group serving as input for the logistic model, outputting the missing probability for the other set of columns. The difference is that after determining the missing probability of the second set, we apply MCAR to the input columns (the first set). Hence, missing values from the second set will depend on the masked values of the first set.
The code for generating masks according to the three missing mechanisms is also provided at https://anonymous.4open.science/r/ICLR-DiffPuter.
D.4 Implementations and hyperparameters
We use a fixed set of hyperparameters, which will save significant efforts in hyperparameter-tuning when applying DiffPuter to more datasets. For the diffusion model, we set the maximum time $T=80$, the noise level $\sigma(t)=t$, which is linear to $t$. The score/denoising neural network $\bm{\epsilon}(\mathbf{x}_{t},t)$ is implemented as a $5$-layer MLP with hidden dimension $1024$. $t$ is transformed to sinusoidal timestep embeddings and then added to $\mathbf{x}_{t}$, which is subsequently passed to the denoising function. When using the learned diffusion model for imputation, we set the number of discrete steps $M=50$ and the number of sampling times per data sample $N=10$. DiffPuter is implemented with Pytorch, and optimized using Adam (Kingma & Ba, 2015) optimizer with a learning rate of $1\times 10^{-4}$.
@@ D.4 Implementations and hyperparameters
We use the same architecture of denoising neural network as in two recent tabular diffusion models for tabular data synthesis (Kotelnikov et al., 2023; Zhang et al., 2024). The denoising MLP takes the current time step $t$ and the feature vector $\mathbf{x}_{t}\in{\mathbb{R}}^{1\times d}$ as input. First, $\mathbf{x}_{t}$ is fed into a linear projection layer that converts the vector dimension to be $d_{\rm hidden}=1024$:
@@
Then, following the practice in TabDDPM (Kotelnikov et al., 2023), the sinusoidal timestep embeddings $\mathbf{t}_{\rm emb}\in{\mathbb{R}}^{1\times d_{\rm hidden}}$ is added to $\mathbf{h}_{0}$ to obtain the input vector $\mathbf{h}_{\rm hidden}$: